# Anytime-valid, Bayes-assisted, Prediction-Powered Inference

### v1

@@ 7 Discussion
where $\varepsilon_{t}=o\left(\frac{1}{\sqrt{t\log t}}\right)$ and $\overline{W}_{t}=\frac{1}{t}\sum_{i=1}^{t}W_{i}$. As shown in Section S5.2, the sequence of balls $\mathcal{C}^{\mathrm{\mathtt{m}BA}}_{\alpha,t}(\overline{W}_{t},\Sigma;\pi)=\mathcal{C}^{\mathrm{\mathtt{m}BA}}_{\alpha,t}(\widehat{\mu}_{t}-\epsilon_{t},\Sigma;\pi)$ forms an exact $(1-\alpha)$ CS for $\mu$. Let
$R_{t}^{\star}$ $=\frac{\sqrt{\Lambda_{\max}(\Sigma)}}{\sqrt{t}}\sqrt{\log\left(\frac{t}{2\pi\alpha^{2}\eta_{t}(\Sigma^{-1/2}(\widehat{\mu}_{t}-\epsilon_{t}))^{2}}\right)}+\|\epsilon_{t}\|.$
As $B\left(\widehat{\mu}_{t},R_{t}^{\ast}\right)\supseteq\mathcal{C}^{\mathrm{\mathtt{m}BA}}_{\alpha,t}(\widehat{\mu}_{t}-\epsilon_{t},\Sigma;\pi)$, the sequence of random balls $B\left(\widehat{\mu}_{t},R_{t}^{\ast}\right)$ also forms an exact $(1-\alpha)$ CS for $\mu$. Define $\mathcal{C}^{\mathrm{\mathtt{m}BA}}_{\alpha,t}(\widehat{\mu}_{t},\widehat{\Sigma}_{t};\pi)=B\left(\widehat{\mu}_{t},R_{t}\right)$, where
$R_{t}=\frac{\sqrt{\Lambda_{\max}(\widehat{\Sigma}_{t})}}{\sqrt{t}}\sqrt{\log\left(\frac{t}{2\pi\alpha^{2}\eta_{t}(\widehat{\Sigma}_{t}^{-1/2}\widehat{\mu}_{t})^{2}}\right)}.$
The consistence of $\widehat{\Sigma}_{t}$ implies that $\|\widehat{\Sigma}_{t}-\Sigma\|=o(1)$ almost surely. By the Courant-Fischer theorem, we have,
$|\Lambda_{\max}\left(\widehat{\Sigma}_{t}\right)-\Lambda_{\max}(\Sigma)|\leq\|\widehat{\Sigma}_{t}-\Sigma\|,$
hence $\Lambda_{\max}\left(\widehat{\Sigma}_{t}\right)\rightarrow\Lambda_{\max}(\Sigma)$ almost surely. So, by Lemma S1, we have
$\eta_{t}(\widehat{\Sigma}_{t}^{-1/2}\widehat{\mu}_{t})\rightarrow\pi\left(\Sigma^{-1/2}\mu\right)\text{\leavevmode\nobreak\ and\leavevmode\nobreak\ }\eta_{t}(\Sigma^{-1/2}(\widehat{\mu}_{t}-\epsilon_{t}))\rightarrow\pi\left(\Sigma^{-1/2}\mu\right)\text{\leavevmode\nobreak\ a.s. as\leavevmode\nobreak\ }t\rightarrow\infty.$
It follows, as in Theorem S3, than $|R_{t}-R^{\star}_{t}|=O(1/\sqrt{t\log t})$ and so $\mathcal{C}^{\mathrm{\mathtt{m}BA}}_{\alpha,t}(\widehat{\mu}_{t},\widehat{\Sigma}_{t};\pi)$ is a $(1-\alpha)$–AsympCS with approximation rate $1/\sqrt{t\log{t}}$. ∎
@@
(Multivariate version of Theorem 3.5) Let $(Y_{t})_{t\geq 1}$ be a sequence of i.i.d. random vectors in $\mathbb{R}^{d}$ with mean $\mu$ and such that $\mathbb{E}\|Y_{1}\|^{2+\delta}<\infty$ for some $\delta>0$. Then, $\mathcal{C}^{\mathrm{\mathtt{m}BA}}_{\alpha,t}(\overline{Y}_{t},\widehat{\Sigma}_{t};\pi)$ is a $(1-\alpha)$–AsympCS with approximation rate $1/\sqrt{t\log{t}}$, where $\overline{Y}_{t}$ is the sample mean and $\widehat{\Sigma}_{t}$ the sample covariance.
Proof.(Theorem S4) By the strong law of large numbers, $\overline{Y}_{t}$ and $\widehat{\Sigma}_{t}$ are consistent estimators of $\mu$ and $\Sigma$, respectively. By the multivariate KMT coupling due to Einmahl (1987) (see Theorem S5), there exists a sequence of i.i.d. Gaussian random vectors $(W_{i})_{i\geq 1}$ with mean $\mu$ and covariance matrix $\Sigma$ such that
$\overline{Y}_{t}=\frac{1}{t}\sum_{i=1}^{t}W_{i}+\varepsilon_{t}\text{\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ where\leavevmode\nobreak\ \leavevmode\nobreak\ \leavevmode\nobreak\ }\varepsilon_{t}=o\left(\frac{1}{t^{1-1/(2+\delta)}}\right)=o\left(\frac{1}{\sqrt{t\log t}}\right).$
The result then follows from Theorem S3. ∎
All other results can be extended to the multivariate case in a similar manner. In the case of Theorem 3.3 and Theorem 3.4, we require a multivariate, unassisted, exact confidence sequence for i.i.d. Gaussian random variables with known variance. For any $\rho>0$, Waudby-Smith et al. (2024a, Equation (29)) propose to use
$\mathcal{C}^{\mathrm{\mathtt{m}NA}}_{\alpha,t}:=\left\{\mu\in\mathbb{R}^{d}:\left\|\overline{Y}_{t}-\mu\right\|<\sqrt{\frac{\Lambda_{\max}(\widehat{\Sigma}_{t})\cdot 9d}{2}\cdot\frac{1+t\rho^{2}}{t^{2}\rho^{2}}\cdot\left[2+\log\left(\frac{\sqrt{1+t\rho^{2}}}{\alpha}\right)\right]}\right\},$
although alternative constructions are possible. To extend our results on control variates and PPI to the multivariate setting, the proofs can be adapted accordingly. In doing so, we will require multivariate versions of the Marcinkiewicz-Zygmund strong law of large numbers (see Ledoux and Talagrand (1991, Theorem 7.9)) and of the law of the iterated logarithm (see Koval (2002, Corollary 1)).
Appendix S6 Mean estimation
@@
The convex loss associated with the estimand $\theta^{*}=\mathbb{E}[Y]$ is the squared loss $\ell_{\theta}(x,y)=(\theta-y)^{2}/2$, whose subgradient with respect to $\theta$ is given by $\ell_{\theta}^{\prime}(x,y)=\theta-y$. As a result of this, the measure of fit $m_{\theta}$ takes the form
@@
As discussed in Section 4, PPI uses the sample mean as an estimator of $m_{\theta}$, which in this case is given by
@@
Given $\widehat{m}_{\theta,n}$ and an estimator $\widehat{\Delta}_{0,n}$ of $\Delta_{0}$, the associated prediction-powered estimator of $\theta^{*}$ is found by solving, in $\theta$, the equation
@@
which match the expressions in Equations 18 and 21, respectively.
As discussed in Section 1, a prediction-powered $(1-\alpha)$ AsympCS $(\mathcal{C}_{\alpha,n}^{\mathrm{avpp}})_{n\geq 1}$ for $\theta$ is defined through Equation 4 by first constructing a valid AsympCS $(\mathcal{C}^{g}_{\alpha,\theta,n})_{n\geq 1}$ for $g_{\theta}$.
Section 5.1 defines valid AsympCS for $g_{\theta}$ that incorporate no prior information. In particular, $\mathcal{C}^{g}_{\alpha,\theta,n}$ is constructed as
$\mathcal{C}^{g}_{\alpha,\theta,n}$ $=\mathcal{C}^{\mathrm{\mathtt{N}A}}_{\alpha,t}(\widehat{g}_{\theta,n},\widehat{\sigma}^{g}_{\theta,n};\rho)$ (S62)
@@
where $\mathcal{C}^{\mathrm{\mathtt{N}A}}_{\alpha,n}$ is defined in Theorem 3.3, $\widehat{g}_{\theta,n}$ is either the PPI estimator $\widehat{g}^{\mathrm{\texttt{PP}}}_{\theta,n}$ or the PPI++ estimator $\widehat{g}^{\mathrm{\texttt{PP+}}}_{\theta,n}$, and $(\widehat{\sigma}^{g}_{\theta,n})^{2}$ is the corresponding variance estimator, as defined in Proposition 5.1. More specifically, under the squared loss, for $\widehat{g}^{\mathrm{\texttt{PP}}}_{\theta,n}$, we have
@@
$\qquad\qquad+\frac{n/N_{n}}{n-1}\sum_{i=1}^{n}\left(Y_{i}-\frac{1}{n}\sum_{k=1}^{n}Y_{i}\right)^{2}.$ (S67)
@@
$\mathcal{C}_{\alpha,n}^{\mathrm{avpp}}$ =$=\left\{\theta\mid 0\in\mathcal{C}^{g}_{\alpha,\theta,n}\right\}$ (S68)
@@
Similarly, Section 5.2 defines valid AsympCS for $g_{\theta}$ that incorporate prior information by means of a zero-mean prior $\pi$ on $\Delta_{\theta}$. In particular, for $\delta\in(0,\alpha)$, Proposition 5.2 first construct a standard $(1-\delta)$ AsympCS $\mathcal{R}_{\delta,\theta,n}$ for $m_{\theta}$, which in the case of mean estimation takes the form
$\mathcal{R}_{\delta,\theta,n}$ $=\mathcal{C}^{\mathrm{\mathtt{N}A}}_{\delta,n}(\widehat{m}_{\theta,n},\widehat{\sigma}^{f}_{\theta,n};\rho)$ (S70)
@@
$=\frac{1}{N_{n}-1}\sum_{j=1}^{N_{n}}\left(f(\widetilde{X}_{i})-\frac{1}{N_{n}}\sum_{k=1}^{N_{n}}f(\widetilde{X}_{k})\right)^{2}.$ (S73)
Next, Proposition 5.3 constructs a Bayes-assisted $(1-(\alpha-\delta))$ AsympCS $\mathcal{T}_{\alpha-\delta,\theta,n}$ for $\Delta_{\theta}$, which under the squared loss takes the form
$\mathcal{T}_{\alpha-\delta,\theta,n}$ $=\mathcal{C}_{\alpha-\delta,n}^{\mathrm{\mathtt{B}A}}(\widehat{\Delta}_{0,n},\widehat{\sigma}^{\Delta}_{\theta,n};\pi)$ (S74)
@@
where $\mathcal{C}^{\mathrm{\mathtt{B}A}}_{\alpha-\delta,n}$ and $\eta_{n}$ are defined in Theorem 3.5, $\widehat{\Delta}_{0,n}$ is either the PPI estimator $\widehat{\Delta}^{\mathrm{\texttt{PP}}}_{0,n}$ or the PPI++ estimator $\widehat{\Delta}^{\mathrm{\texttt{PP}}}_{0,n}$, and $(\widehat{\sigma}^{\Delta}_{\theta,n})^{2}$ is the corresponding variance estimator, as defined in Proposition 5.3. In the case of mean estimation, $(\widehat{\sigma}^{\Delta}_{\theta,n})^{2}$ takes the form
@@
$\mathcal{C}_{\alpha,n}^{\mathrm{avpp}}$ =$=\left\{\theta\mid 0\in\mathcal{C}^{g}_{\alpha,\theta,n}\right\}$ (S81)
@@
which matches the expression in Equation 23, and where $\widehat{\theta}_{n}$ is either the PPI estimator $\widehat{\theta}_{n}^{\mathrm{\texttt{PP}}}$ or the PPI++ estimator $\widehat{\theta}_{n}^{\mathrm{\texttt{PP+}}}$.
@@
S7.1 Implementation
All experiments were run locally on an Intel Core i7-11850H CPU.
S7.2 Datasets
Here we briefly describe each dataset used for the real data experiments in Section 6.2. The flights dataset was downloaded from Kaggle [^1] https://www.kaggle.com/datasets/shubhambathwal/flight-price-prediction, while all the others are available as part of the ppi-py package [^2] https://pypi.org/project/ppi-python/.
@@
For each of $103333$ economy class flight tickets, the flights dataset reports the ticket price, as well as the prediction of a gradient-boosted tree for the same quantity. The goal is to estimate the average price of a flight, i.e. $\theta^{*}=\mathbb{E}[Y]\in\mathbb{R}$.
@@
For each of $1596$ parcels of land in the Amazon rainforest (Bullock et al., 2020), the forest dataset reports whether the parcel has been subject to deforestation ($Y_{i}\in\{0,1\}$), as well as the prediction of a gradient-boosted tree model for the probability of the same event ($f(X_{i})\in[0,1]$). The goal is to estimate the fraction of Amazon rainforest lost to deforestation, i.e. $\theta^{*}=\mathbb{E}[Y]\in[0,1]$.
@@
For each of $16743$ images from the Galaxy Zoo 2 initiative (Willett et al., 2013), the galaxies dataset reports whether the galaxy has spiral arms ($Y_{i}\in\{0,1\}$), as well as the prediction of a ResNet50 model (He et al., 2016) for the probability of the same event ($f(X_{i})\in[0,1]$). The goal is to estimate the fraction of galaxies with spiral arms, i.e. $\theta^{*}=\mathbb{E}[Y]\in[0,1]$.
S7.3 AsympCS hyperparameters
Here we discuss the hyperparameters of the PPI and PPI++ AsympCS procedures defined in Section 5.
@@
The unassisted prediction-powered AsympCS discussed in Section 5.1 requires the specification of the parameter $\rho$ for Equation 5 in Theorem 3.3. As mentioned at the end of Section 3.1, $\rho$ can be chosen so as to minimise the width of the interval at a specified time. In particular, as shown in Waudby-Smith et al. (2024a, Appendix B.2), setting
$\rho=\sqrt{\frac{-W_{-1}(-\alpha^{2}\exp(-1))-1}{t^{*}}},$ (S83)
@@
where $W_{-1}$ is the lower branch of the Lambert $W$ function, minimises the width of the interval at time $t^{*}$.
@@
The Bayes-assisted prediction-powered AsympCS discussed in Section 5.2 requires the specification of both the parameter $\rho$ used for the unassisted AsympCS for the measure of fit $m_{\theta}$ (Proposition 5.2) and the scale parameter $\tau$ of the prior $\pi$ for the Bayes-assisted AsympCS for the rectifier $\Delta_{\theta}$ (Proposition 5.3). While the former can be chosen as above, the same approach does not work for the latter, as the prior scale that minimises the width of the Bayes-assisted interval at a specified time $t$ depends on the observed value of $\overline{Y}_{t}/\widehat{\sigma}_{t}$ in Equation 8 of Theorem 3.5. Instead, we propose the following heuristic for choosing $\tau$. If $Z_{1},Z_{2},\ldots|\mu\overset{iid}{\sim}\mathcal{N}(\mu,1)$, then the posterior mean $\mathbb{E}[\mu|Z_{1},\dots,Z_{t}]$ after $t$ observations under a Gaussian prior $\mu\sim\mathcal{N}(\mu_{0},\tau^{2})$ is given by
@@
where the two terms on the right-hand side measure the influence of the prior and the data on the posterior mean, respectively. Noticing that the Gaussian likelihood leading to the posterior mean (S84) is of the same form as the one implicitly used for the construction of Bayes-assisted AsympCS (see e.g. Equation 7), we choose $\tau$ so that the prior and the data have the same influence on the posterior mean (S84) at time $t^{*}$, i.e.
$\tau=\frac{1}{\sqrt{t^{*}}}.$ (S85)
Appendix S8 reports the hyperparameter value used for each experiment in terms of $t^{*}$. Notice that, as discussed in Section 6, the initial $N_{n}$ is set large enough to rule out any uncertainty on the measure of fit $m_{\theta}$. As a result of this, the only hyperparameter that needs to be chosen for the Bayes-assisted procedure is the prior scale $\tau$.

### v2

@@ 7 Discussion
where $\varepsilon_{t}=o\left(\frac{1}{\sqrt{t\log t}}\right)$ and $\overline{W}_{t}=\frac{1}{t}\sum_{i=1}^{t}W_{i}$. As stated in Section˜S4.2, the sequence of balls $\mathcal{C}^{\mathrm{\tt mBA}}_{\alpha,t}(\overline{W}_{t},\Sigma;\uppi)=\mathcal{C}^{\mathrm{\tt mBA}}_{\alpha,t}(\widehat{\mu}_{t}-\varepsilon_{t},\Sigma;\uppi)$ forms an exact $(1-\alpha)$ CS for $\mu$. Let
$R_{t}^{\star}$ $=\frac{\sqrt{\Lambda_{\max}(\Sigma)}}{\sqrt{t}}\sqrt{\log\left(\frac{t^{d}}{(2\pi)^{d}\alpha^{2}\eta_{t}(\Sigma^{-1/2}(\widehat{\mu}_{t}-\varepsilon_{t}))^{2}}\right)}+\|\varepsilon_{t}\|.$
As $B\left(\widehat{\mu}_{t},R_{t}^{\star}\right)\supseteq\mathcal{C}^{\mathrm{\tt mBA}}_{\alpha,t}(\widehat{\mu}_{t}-\varepsilon_{t},\Sigma;\uppi)$, the sequence of random balls $B\left(\widehat{\mu}_{t},R_{t}^{\star}\right)$ also forms an exact $(1-\alpha)$ CS for $\mu$. Define $\mathcal{C}^{\mathrm{\tt mBA}}_{\alpha,t}(\widehat{\mu}_{t},\widehat{\Sigma}_{t};\uppi)=B\left(\widehat{\mu}_{t},R_{t}\right)$, where
$R_{t}=\frac{\sqrt{\Lambda_{\max}(\widehat{\Sigma}_{t})}}{\sqrt{t}}\sqrt{\log\left(\frac{t^{d}}{(2\pi)^{d}\alpha^{2}\eta_{t}(\widehat{\Sigma}_{t}^{-1/2}\widehat{\mu}_{t})^{2}}\right)}.$
By the Courant-Fischer theorem, we have,
$|\Lambda_{\max}\left(\widehat{\Sigma}_{t}\right)-\Lambda_{\max}(\Sigma)|\leq\|\widehat{\Sigma}_{t}-\Sigma\|=o(1/\log t)\text{ a.s.}$
Additionally, by Lemma˜S2, we have
$\eta_{t}(\widehat{\Sigma}_{t}^{-1/2}\widehat{\mu}_{t})\rightarrow\uppi\left(\Sigma^{-1/2}\mu\right)\text{~and~}\eta_{t}(\Sigma^{-1/2}(\widehat{\mu}_{t}-\varepsilon_{t}))\rightarrow\uppi\left(\Sigma^{-1/2}\mu\right)\text{~a.s. as~}t\rightarrow\infty.$
It follows that
$R_{t}-(R_{t}^{\star}-\|\varepsilon_{t}\|)$ $=\frac{R_{t}^{2}-(R_{t}^{\star}-\|\varepsilon_{t}\|)^{2}}{R_{t}+(R_{t}^{\star}-\|\varepsilon_{t}\|)}$
$\sim\frac{(\Lambda_{\max}(\widehat{\Sigma}_{t})-\Lambda_{\max}(\Sigma))\frac{\log t^{d}}{t}}{\sqrt{\Lambda_{\max}(\Sigma)\frac{\log t^{d}}{t}}}$
$=o(1/\sqrt{t\log t})$
a.s. Then $|R_{t}-R^{\star}_{t}|=o(1/\sqrt{t\log t})$ and so $\mathcal{C}^{\mathrm{\tt mBA}}_{\alpha,t}(\widehat{\mu}_{t},\widehat{\Sigma}_{t};\uppi)$ is a $(1-\alpha)$–AsympCS with approximation rate $1/\sqrt{t\log{t}}$. ∎
@@
(Multivariate version of Theorem˜3) Let $(Y_{t})_{t\geq 1}$ be a sequence of i.i.d. random vectors in $\mathbb{R}^{d}$ with mean $\mu$ and such that $\mathbb{E}\|Y_{1}\|^{2+\delta}<\infty$ for some $\delta>0$. Then, $\mathcal{C}^{\mathrm{\tt mBA}}_{\alpha,t}(\overline{Y}_{t},\widehat{\Sigma}_{t};\uppi)$ is a $(1-\alpha)$–AsympCS with approximation rate $1/\sqrt{t\log{t}}$, where $\overline{Y}_{t}$ is the sample mean and $\widehat{\Sigma}_{t}$ the sample covariance.
Proof.
(Theorem˜S13) By the strong law of large numbers, $\overline{Y}_{t}$ and $\widehat{\Sigma}_{t}$ are consistent estimators of $\mu$ and $\Sigma$, respectively. By the multivariate KMT coupling due to Einmahl [5] (see Theorem˜S9), there exists a sequence of i.i.d. Gaussian random vectors $(W_{i})_{i\geq 1}$ with mean $\mu$ and covariance matrix $\Sigma$ such that
$\overline{Y}_{t}=\frac{1}{t}\sum_{i=1}^{t}W_{i}+\varepsilon_{t}\text{~~~where~~~}\varepsilon_{t}=o\left(\frac{1}{t^{1-1/(2+\delta)}}\right)=o\left(\frac{1}{\sqrt{t\log t}}\right).$
The result then follows from Theorem˜S12. ∎
All other results can be extended to the multivariate case in a similar manner. In the case of Theorem˜1 and Theorem˜2, we require a multivariate, non-assisted, exact confidence sequence for i.i.d. Gaussian random variables with known variance. For any $\rho>0$, Waudby-Smith et al. [6, Equation (29)] propose to use
$\mathcal{C}^{\mathrm{\tt mNA}}_{\alpha,t}:=\left\{\mu\in\mathbb{R}^{d}:\left\|\overline{Y}_{t}-\mu\right\|<\sqrt{\frac{\Lambda_{\max}(\widehat{\Sigma}_{t})\cdot 9d}{2}\cdot\frac{1+t\rho^{2}}{t^{2}\rho^{2}}\cdot\left[2+\log\left(\frac{\sqrt{1+t\rho^{2}}}{\alpha}\right)\right]}\right\},$
although alternative constructions are possible. To extend our results on control variates and PPI to the multivariate setting, the proofs can be adapted accordingly. In doing so, we will require multivariate versions of the Marcinkiewicz-Zygmund strong law of large numbers (see Ledoux and Talagrand [22, Theorem 7.9]) and of the law of the iterated logarithm (see Koval [23, Corollary 1]).
Appendix S5 Derivations for prediction-powered mean estimation
@@
The convex loss associated with the estimand $\theta^{\star}=\mathbb{E}[Y]$ is the squared loss $\ell_{\theta}(x,y)=(\theta-y)^{2}/2$, whose subgradient with respect to $\theta$ is given by $\ell_{\theta}^{\prime}(x,y)=\theta-y$. As a result of this, the measure of fit $m_{\theta}$ takes the form
@@
As discussed in Section˜4, PPI uses the sample mean as an estimator of $m_{\theta}$, which in this case is given by
@@
Given $\widehat{m}_{\theta,n}$ and an estimator $\widehat{\Delta}_{0,n}$ of $\Delta_{0}$, the associated prediction-powered estimator of $\theta^{\star}$ is found by solving, in $\theta$, the equation
@@
which match the expressions in Equations˜22 and 25, respectively.
As discussed in Section˜1, a prediction-powered $(1-\alpha)$ AsympCS $(\mathcal{C}_{\alpha,n}^{\mathrm{avpp}})_{n\geq 1}$ for $\theta$ is defined through Equation˜4 by first constructing a valid AsympCS $(\mathcal{C}^{g}_{\alpha,\theta,n})_{n\geq 1}$ for $g_{\theta}$.
Section˜5.1 defines valid AsympCS for $g_{\theta}$ that incorporate no prior information. In particular, $\mathcal{C}^{g}_{\alpha,\theta,n}$ is constructed as
$\mathcal{C}^{g}_{\alpha,\theta,n}$ $=\mathcal{C}^{\mathrm{\tt NA}}_{\alpha,t}(\widehat{g}_{\theta,n},\widehat{\sigma}^{g}_{\theta,n};\rho)$ (S66)
@@
where $\mathcal{C}^{\mathrm{\tt NA}}_{\alpha,n}$ is defined in Theorem˜1, $\widehat{g}_{\theta,n}$ is either the PPI estimator $\widehat{g}^{\mathrm{\texttt{PP}}}_{\theta,n}$ or the PPI++ estimator $\widehat{g}^{\mathrm{\texttt{PP+}}}_{\theta,n}$, and $(\widehat{\sigma}^{g}_{\theta,n})^{2}$ is the corresponding variance estimator, as defined in Proposition˜3. More specifically, under the squared loss, for $\widehat{g}^{\mathrm{\texttt{PP}}}_{\theta,n}$, we have
@@
$\qquad\qquad+\frac{n/N_{n}}{n-1}\sum_{i=1}^{n}\left(Y_{i}-\frac{1}{n}\sum_{k=1}^{n}Y_{k}\right)^{2}.$ (S71)
@@
$\mathcal{C}_{\alpha,n}^{\mathrm{avpp}}$ $=\left\{\theta\mid 0\in\mathcal{C}^{g}_{\alpha,\theta,n}\right\}$ (S72)
@@
Similarly, Section˜5.2 defines valid AsympCS for $g_{\theta}$ that incorporate prior information by means of a zero-mean prior $\pi$ on $\Delta_{\theta}$. In particular, for $\delta\in(0,\alpha)$, Proposition˜4 first construct a standard $(1-\delta)$ AsympCS $\mathcal{R}_{\delta,\theta,n}$ for $m_{\theta}$, which in the case of mean estimation takes the form
$\mathcal{R}_{\delta,\theta,n}$ $=\mathcal{C}^{\mathrm{\tt NA}}_{\delta,n}(\widehat{m}_{\theta,n},\widehat{\sigma}^{f}_{\theta,n};\rho)$ (S74)
@@
$=\frac{1}{N_{n}-1}\sum_{j=1}^{N_{n}}\left(f(\widetilde{X}_{j})-\frac{1}{N_{n}}\sum_{k=1}^{N_{n}}f(\widetilde{X}_{k})\right)^{2}.$ (S77)
Next, Proposition˜5 constructs a Bayes-assisted $(1-(\alpha-\delta))$ AsympCS $\mathcal{T}_{\alpha-\delta,\theta,n}$ for $\Delta_{\theta}$, which under the squared loss takes the form
$\mathcal{T}_{\alpha-\delta,\theta,n}$ $=\mathcal{C}_{\alpha-\delta,n}^{\mathrm{\tt BA}}(\widehat{\Delta}_{0,n},\widehat{\sigma}^{\Delta}_{\theta,n};\pi)$ (S78)
@@
where $\mathcal{C}^{\mathrm{\tt BA}}_{\alpha-\delta,n}$ and $\eta_{n}$ are defined in Theorem˜3, $\widehat{\Delta}_{0,n}$ is either the PPI estimator $\widehat{\Delta}^{\mathrm{\texttt{PP}}}_{0,n}$ or the PPI++ estimator $\widehat{\Delta}^{\mathrm{\texttt{PP+}}}_{0,n}$, and $(\widehat{\sigma}^{\Delta}_{\theta,n})^{2}$ is the corresponding variance estimator, as defined in Proposition˜5. In the case of mean estimation, $(\widehat{\sigma}^{\Delta}_{\theta,n})^{2}$ takes the form
@@
$\mathcal{C}_{\alpha,n}^{\mathrm{avpp}}$ $=\left\{\theta\mid 0\in\mathcal{C}^{g}_{\alpha,\theta,n}\right\}$ (S85)
@@
which matches the expression in Equation˜27, and where $\widehat{\theta}_{n}$ is either the PPI estimator $\widehat{\theta}_{n}^{\mathrm{\texttt{PP}}}$ or the PPI++ estimator $\widehat{\theta}_{n}^{\mathrm{\texttt{PP+}}}$.
@@
S6.1 Implementation
Code implementing our method is written in Python and made available at https://github.com/stefanocortinovis/ppi-cs. All experiments were run locally on an Apple Silicon M4 Pro CPU with 24GB of memory.
S6.2 Datasets
Here we briefly describe each dataset used for the real data experiments in Section˜6.2. The flights dataset was downloaded from Kaggle [^1] https://www.kaggle.com/datasets/shubhambathwal/flight-price-prediction, while all the others are available as part of the ppi-python package [^2] https://pypi.org/project/ppi-python/.
@@
For each of $103333$ economy class flight tickets, the flights dataset reports the ticket price ($Y_{i}\in\mathbb{R}$), as well as the prediction of a gradient-boosted tree for $Y_{i}$ ($f(X_{i})\in\mathbb{R}$). The goal is to estimate the average price of a flight, i.e. $\theta^{\star}=\mathbb{E}[Y]\in\mathbb{R}$.
@@
For each of $1596$ parcels of land in the Amazon rainforest [24], the forest dataset reports whether the parcel has been subject to deforestation ($Y_{i}\in\{0,1\}$), as well as the prediction of a gradient-boosted tree model for the probability of $Y_{i}$ being equal to one ($f(X_{i})\in[0,1]$). The goal is to estimate the fraction of Amazon rainforest lost to deforestation, i.e. $\theta^{\star}=\mathbb{E}[Y]\in[0,1]$.
@@
For each of $16743$ images from the Galaxy Zoo 2 initiative [25], the galaxies dataset reports whether the galaxy has spiral arms ($Y_{i}\in\{0,1\}$), as well as the prediction of a ResNet50 model [26] for the probability of $Y_{i}$ being equal to one ($f(X_{i})\in[0,1]$). The goal is to estimate the fraction of galaxies with spiral arms, i.e. $\theta^{\star}=\mathbb{E}[Y]\in[0,1]$.
Census.
For each of $380091$ individuals from the 2019 California census, the census dataset reports the individual’s age ($X_{i}\in\mathbb{R}$) and yearly income ($Y_{i}\in\mathbb{R}$), as well as the prediction of a gradient-boosted tree model trained on the previous year’s data for $Y_{i}$ ($f(X_{i})\in\mathbb{R}$). The goal is to estimate the ordinary least squares (OLS) regression coefficient when regressing income on age. We preprocess the data by excluding non-positive incomes ($Y_{i}\leq 0$ or $f(X_{i})\leq 0$), and applying a log-transformation to both the response $Y_{i}$ and the prediction $f(X_{i})$. This results in a dataset of $268118$ individuals.
Healthcare.
For each of $318215$ individuals from the 2019 California census, the healthcare dataset reports the individual’s yearly income ($X_{i}\in\mathbb{R}$) and whether they have health insurance ($Y_{i}\in\{0,1\}$), as well as the prediction of a gradient-boosted tree model trained on the previous year’s data for the probability of $Y_{i}$ being equal to one ($f(X_{i})\in[0,1]$). The goal is to estimate the logistic regression coefficient when regressing health insurance status on income. As above, we preprocess the data by excluding non-positive incomes ($X_{i}\leq 0$), and applying a log-transformation to the covariate $X_{i}$. This results in a dataset of $270214$ individuals.
Genes.
For each of $61150$ gene promoter sequences [27], the genes dataset reports the expression level of the gene induced by the promoter ($Y_{i}\in\mathbb{R}$), as well as the prediction of a transformer model for $Y_{i}$ ($f(X_{i})\in\mathbb{R}$). The goal is to estimate the median expression level across sequences. We preprocess the data by applying a log-transformation to the response $Y_{i}$.
S6.3 Predictor performance
Table˜S1 reports the performance of the predictors used for each real data dataset above, measured in terms of normalised root mean squared error (NRMSE) for regression (R) tasks (flights, census, genes) and cross-entropy (CE) for the binary classification (C) tasks (forest, galaxies, healthcare).
Table S1: Predictor performance on real data datasets.
Dataset Flights Forest Galaxies Census Healthcare Genes
Task R C C R C R
Performance $0.20$ $0.31$ $0.29$ $0.11$ $0.36$ $0.33$
While we report these for completeness, we emphasise that non-assisted PPI improves over classical inference in the presence of correlation between the predictions and the true labels, regardless of the absolute predictive performance (see e.g. Figure˜2). On the other hand, while knowledge on the predictive performance can be used to choose a suitable prior for Bayes-assisted PPI, the latter is placed on the rectifier $\Delta_{\theta}$, which depends on the downstream inference task, thereby making the relationship between predictive performance and efficiency gains less direct.
S6.4 AsympCS hyperparameters
Here we discuss the hyperparameters of the PPI and PPI++ AsympCS procedures defined in Section˜5.
@@
The non-assisted prediction-powered AsympCS discussed in Section˜5.1 requires the specification of the parameter $\rho$ for Equation˜6 in Theorem˜1. As mentioned at the end of Section˜3.1, $\rho$ can be chosen so as to minimise the width of the interval at a specified time. In particular, as shown in Waudby-Smith et al. [6, Appendix B.2], setting
$\rho=\sqrt{\frac{-W_{-1}(-\alpha^{2}\exp(-1))-1}{t^{\star}}},$ (S87)
@@
where $W_{-1}$ is the lower branch of the Lambert $W$ function, minimises the width of the interval at time $t^{\star}$.
@@
The Bayes-assisted prediction-powered AsympCS discussed in Section˜5.2 requires the specification of both the parameter $\rho$ used for the non-assisted AsympCS for the measure of fit $m_{\theta}$ (Proposition˜4) and the scale parameter $\tau$ of the prior $\pi$ for the Bayes-assisted AsympCS for the rectifier $\Delta_{\theta}$ (Proposition˜5). While the former can be chosen as above, the same approach does not work for the latter, as the prior scale that minimises the width of the Bayes-assisted interval at a specified time $t$ depends on the observed value of $\overline{Y}_{t}/\widehat{\sigma}_{t}$ in Equation˜9 of Theorem˜3. Instead, we propose the following heuristic for choosing $\tau$. If $Z_{1},Z_{2},\ldots|\mu\overset{\mathrm{iid}}{\sim}\mathcal{N}(\mu,1)$, then the posterior mean $\mathbb{E}[\mu|Z_{1},\dots,Z_{t}]$ after $t$ observations under a Gaussian prior $\mu\sim\mathcal{N}(\mu_{0},\tau^{2})$ is given by
@@
where the two terms on the right-hand side measure the influence of the prior and the data on the posterior mean, respectively. Noticing that the Gaussian likelihood leading to the posterior mean (S88) is of the same form as the one implicitly used for the construction of Bayes-assisted AsympCS (see e.g. Equation˜8), we choose $\tau$ so that the prior and the data have the same influence on the posterior mean (S88) at time $t^{\star}$, i.e.
$\tau=\frac{1}{\sqrt{t^{\star}}}.$ (S89)
Appendix˜S7 reports the hyperparameter value used for each experiment in terms of $t^{\star}$. Notice that, as discussed in Section˜6, the initial $N_{n}$ is set large enough to rule out any uncertainty on the measure of fit $m_{\theta}$. As a result of this, the only hyperparameter that needs to be chosen for the Bayes-assisted procedure is the prior scale $\tau$.